Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by akhpad » Mon May 17, 2010 1:21 am
Bill has a small deck of 12 playing cards made up of only 2 suits of 6 cards each. Each of the 6 cards within a suit has a different value from 1 to 6; thus, for each value from 1 to 6, there are two cards in the deck with that value. Bill likes to play a game in which he shuffles the deck, turns over 4 cards, and looks for pairs of cards that have the same value. What is the chance that Bill finds at least one pair of cards that have the same value?

Solution:
(1, 1), (2, 2), (3, 3), (4, 4), (5, 5), (6, 6)

Method 1:
First find out chances of getting all cards different
Probability of selecting 1st card = 12/12
Probability of selecting 2st card = 10/11
Probability of selecting 3st card = 8/10
Probability of selecting 4st card = 6/9

Probability of selecting
Probability of selecting (all 4 cards different) = (12/12) * (10/11) * (8/10) * (6/9) = 16/33
Probability of selecting (at least one pain) = 1 – 16/33 = 17/33

Method 2:
Total combination with one or more pair = 6 * (10C2) – 6C2 = 5*51
Probability of selecting at least one pain = 5*51 / 495 = 17/33

Method 3:
Total combination with all different = 12 * 10 * 8 * 6 / 4! =240
Total no of selections = 12C4 = 495
Probability of selecting at least one pain = (495 – 240) / 495 = 17/33

Hence required probability = 17/33

My Q: Why do we need to divide by 4! in Method 3?
Join the discussion
Source: — Problem Solving |

by sanju09 » Mon May 17, 2010 3:56 am
akhp77 wrote:Bill has a small deck of 12 playing cards made up of only 2 suits of 6 cards each. Each of the 6 cards within a suit has a different value from 1 to 6; thus, for each value from 1 to 6, there are two cards in the deck with that value. Bill likes to play a game in which he shuffles the deck, turns over 4 cards, and looks for pairs of cards that have the same value. What is the chance that Bill finds at least one pair of cards that have the same value?

Solution:
(1, 1), (2, 2), (3, 3), (4, 4), (5, 5), (6, 6)

Method 1:
First find out chances of getting all cards different
Probability of selecting 1st card = 12/12
Probability of selecting 2st card = 10/11
Probability of selecting 3st card = 8/10
Probability of selecting 4st card = 6/9

Probability of selecting
Probability of selecting (all 4 cards different) = (12/12) * (10/11) * (8/10) * (6/9) = 16/33
Probability of selecting (at least one pain) = 1 – 16/33 = 17/33

Method 2:
Total combination with one or more pair = 6 * (10C2) – 6C2 = 5*51
Probability of selecting at least one pain = 5*51 / 495 = 17/33

Method 3:
Total combination with all different = 12 * 10 * 8 * 6 / 4! =240
Total no of selections = 12C4 = 495
Probability of selecting at least one pain = (495 – 240) / 495 = 17/33

Hence required probability = 17/33

My Q: Why do we need to divide by 4! in Method 3?
You have shown a great piece of work, Akhilesh.

Well, in your method 3, 12 * 10 * 8 * 6 shows all possible arrangements of the four chosen cards, whereas we are looking for total combinations of those four cards, and that nCr = nPr/r!, therefore we need to divide 12 * 10 * 8 * 6 by 4! for the correct answer to the question.
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion

by harshavardhanc » Mon May 17, 2010 4:09 am
sanju09 wrote:
Well, in your method 3, 12 * 10 * 8 * 6 shows all possible arrangements of the four chosen cards, whereas we are looking for total combinations of those four cards, and that nCr = nPr/r!, therefore we need to divide 12 * 10 * 8 * 6 by 4! for the correct answer to the question.
did you mean 12*10*8*6 is some nPr ? How?
Regards,
Harsha
Join the discussion

by sanju09 » Mon May 17, 2010 4:33 am
harshavardhanc wrote:
sanju09 wrote:
Well, in your method 3, 12 * 10 * 8 * 6 shows all possible arrangements of the four chosen cards, whereas we are looking for total combinations of those four cards, and that nCr = nPr/r!, therefore we need to divide 12 * 10 * 8 * 6 by 4! for the correct answer to the question.
did you mean 12*10*8*6 is some nPr ? How?
sort of

there are four place numbers being filled in 12, 10, 8, and 6 number of ways in any order. Hence, all arrangements were considered here, we need to divide it by 4! in order to get the favorite combinations only.
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion

by akhpad » Tue May 18, 2010 5:00 am
Your answer to my question looks good. Thanks

Some more comment from other people please.
Join the discussion